# On the Potential of an Independent Avatar to Augment Metaverse User Socialization Time

### v1

@@ 3.3. Social presence
In the social psychology literature, our notion of “independent avatar” would naturally fall in the category of “agent”. Research works which have investigated the impact of the (perceived) agency on social presence typically assume the user as an actual person or a computerized character prior to the interaction. As the survey (Oh et al., 2018) nicely points out, the majority of the related literature converges to the conclusion that avatars generally elicit greater social presence value than agents, in the sense that users typically feel higher levels of social presence when a virtual counterpart was thought to be controlled by an actual person rather than by a computer program.
@@
(8) $y_{v}=f_{s}(\tilde{x}_{v})=\beta_{v}\cdot\tilde{x}_{v}$
time, with, according to (Kiyokawa et al., 2012),
(9) $\mathbf{E}[\beta_{v}]=\frac{4.182}{3.236}=1.29,$
where $\tilde{x}_{v}$ would be the time that $u$ would need to spend socializing with $v$ instead.
3.4. Social cues and compression
@@
(11) $z_{v}=c\cdot\delta\cdot y_{v}=\gamma\cdot y_{v},$
@@
According to known works on compression ratios, such as (Goldstein et al., 1999) and(Ghalandari et al., 2022), we can set
(12) $\mathbb{E}[c]=0.54.$
Also, social presence can be measured in several directions (verbal, visual, auditory, etc.). (Chen et al., 2023) provides a quantitative survey-based assessment that captures both the visual and verbal efficiency of chatbots, using a 7-point Likert scale. As an example, in the best case, highly anthropomorphic chatbots with a human-like conversation style reach a score of 6 out of 7. We use these results to assign a value to $\delta$. Specifically, in the above example, we say that
(13) $\mathbb{E}[\delta]=7/6.$
@@
4. Maximizing the user spare time
@@ 4.1. Problem formulation
For a user $u$ and independent avatar $a$, given a set of contacts $V$ and their respective initial non-avatar-mediated time partitioning $\tilde{x_{1}}+...+\tilde{x_{n}}=\tilde{X}$, maximize the spare time of $u$ (i.e., $\tilde{X}-X-Z$) by also using $a$’s avatar-mediated time for socialization in a resulting total combined time partitioning $(x_{1}+y_{1})+...+(x_{n}+y_{n})=X+Y$.
@@
(14) $\min\quad$ $\sum_{v}x_{v}+\gamma\cdot y_{v}$
@@
(15) $x_{v}+\frac{1}{\beta_{v}}\cdot y_{v}\geq\tilde{x}_{v}$ $\forall v$
(16) $\sum_{v}x_{v}+\gamma\cdot y_{v}\leq\tilde{X}$
(17) $\sum_{v}y_{v}\leq Y$
(18) $\sum_{v}\gamma\cdot y_{v}\leq Z$
(19) $x_{v},y_{v}\geq 0$ $\forall v$
The objective function (Eq. 14) minimizes the non-avatar-mediated time and the debriefing time, thus maximizing the available non-avatar-mediated spare time on the user side. Constraints in Eq. 15 guarantee that the resulting effective non-avatar-mediated and avatar-mediated socializing time for each $v$ is not less than the initial non-avatar-mediated time. Constraint in Eq. 16 guarantees that the non-avatar-mediated time and the debriefing time do not surpass the total available time of the user. Constraint in Eq. 17 guarantees that the avatar-mediated time does not surpass the total available time of the avatar. Constraint 18 guarantees that the debriefing process abides by the non-distractibility rule introduced with Eq. 6. Finally, constraints in Eq. 18 guarantee that time values are zero or positive.
4.2. Problem feasibility
@@
(20) $x_{v}=\tilde{x}_{v}-\frac{1}{\beta}y_{v}.$
Effectively, we just need to find a solution for $y_{v}$ and the values for $x_{v}$ will follow. This allows us to rewrite the optimization problem as follows.
(21) $\min\quad$ $\tilde{X}+\left(\gamma-\frac{1}{\beta}\right)\sum_{v}y_{v}$
@@
(22) $\left(\gamma-\frac{1}{\beta}\right)\sum_{v}y_{v}\leq 0$
(23) $\sum_{v}y_{v}\leq\min\left\{Y,\frac{Z}{\gamma}\right\}$
(24) $0\leq y_{v}\leq\beta\tilde{x}_{v}$ $\forall v$
Let us now look at the first constraint (Eq. 22). Since, according to the constraint in Eq. 24, $y_{v}$ are always positive, the first constraint can only be satisfied if $\gamma\leq\frac{1}{\beta}$ (case \⃝raisebox{-0.9pt}{A}) or if $y_{v}=0$ for all $v$ (case \⃝raisebox{-0.9pt}{B}). The discussion in Section 3 tells us that in a realistic setting $\frac{1}{\beta}\sim 0.78$ and $\gamma\sim 0.63$, so we are operating in case \⃝raisebox{-0.9pt}{A} and non-trivial solutions (i.e., $y_{v}>0$) are possible. The remaining constraint (Eq. 23) is a simple cap on the total avatar time. Its concrete meaning is that the total time the avatar spends socializing cannot be greater than the overall time in the avatar’s capacity for socialization or than the maximum debriefing capacity of the user.
Figure 1. Contour plot of the objective function in Eq. 21 when the socialization involves two alters. The dashed blue line corresponds to the constraint in Eq. 23, the dashed green lines to those in Eq. 24. The red line is the minimizing solution for case \⃝raisebox{-0.9pt}{A}, the red point for case \⃝raisebox{-0.9pt}{B}. The plot is obtained with $\tilde{x}_{1}=40,\tilde{x}_{2}=15,\gamma=0.63,\frac{1}{\beta}=0.78,Y=45,Z=45$.
In light of the reformulated problem, we can now revisit the objective function $\min\tilde{X}+(\gamma-\frac{1}{\beta})\sum_{v}y_{v}$. Let us distinguish the two cases, \⃝raisebox{-0.9pt}{A} and \⃝raisebox{-0.9pt}{B}, that we have already introduced. When $(\gamma-\frac{1}{\beta})>0$⃝raisebox{-0.9pt}{B}, the objective function is minimized when $y_{v}=0$ for all $v$. Vice versa, when $(\gamma-\frac{1}{\beta})\leq 0$⃝raisebox{-0.9pt}{A}, as in the realistic conditions discussed in Section 3, the objective function is minimized when $y_{v}$ take their highest values within the feasible region. As illustrated in Figure 1, in case \⃝raisebox{-0.9pt}{A}, the objective function is minimized when $y_{v}$ lies on the red solid line, while in case \⃝raisebox{-0.9pt}{B} when $y_{v}=0$.
@@
5. Preliminary insights
We wrapped the previous section with a qualitative finding: when the debriefing process is more efficient than the avatar socialization process, using the avatar is always convenient. In this section, we delve into a quantitative analysis of this benefit. To this aim, we leverage the model provided in (Conti et al., 2012) for generating realistic ego networks. We fix $\beta$ to $1.29$ in our optimization problem and we vary $\gamma$ between $[0,\frac{1}{\beta})$ (i.e., we remain in the range where case \⃝raisebox{-0.9pt}{A} holds but we assume, as it is realistic, that there will be some variability in the $\gamma$ values). We test scenarios where the total avatar time $Y$ is greater, equal, or smaller than the total user socialization time without the avatar ($\tilde{X}$). In Figure 2, we plot the amount of spare time gained by the user by leveraging the avatar. When $\gamma$ and $\frac{1}{\beta}$ are close (i.e., their difference is near zero), the gained spare time is very small. However, as the debriefing becomes more and more efficient with respect to the avatar socialization, then the time saved by the user becomes increasingly higher.
@@ 5. Preliminary insights
To quantify the potential impact of the spare time, we compare the saved time against the time an average user spends on individual alters of its ego network layers. We extract these numbers from (Conti et al., 2012), where statistics for three concentric layers are reported. The layers are referred to as the “support clique”, “sympathy group”, and “active network”, with average cumulative sizes of 4.6, 14.3, and 132.5 alters, respectively. This structure deviates from the standard ego network structure due to the exclusion of the “affinity group” layer, whose properties are not well-defined in the related literature. From (Conti et al., 2012), we obtain $t_{\textrm{active}}=8.81h$, $t_{\textrm{sympathy}}=38.72h$, $t_{\textrm{support}}=74h$. Considering together the time spent on all the alters in each layer, we obtain $T_{\textrm{active}}=1041.51h$, $T_{\textrm{sympathy}}=375.83h$, $T_{\textrm{support}}=340.65h$. It is evident that when the amount of time saved is high enough, users might have the opportunity to reinvest it in expanding their ego network circles and enhancing their social connections. For instance, with approximately 1000 spare hours, one could potentially double the active layer of their ego network, or with around 350 spare hours, strengthen their support clique. While fully characterizing the dynamics of this augmentation is beyond the scope of this work, and the relationship between time saved and new connections gained might not be as direct and straightforward, these initial insights suggest that avatars in the Metaverse have the potential to increase a user’s available time – a resource that could be redirected toward growing their personal social network. In theory, this could result in surpassing Dunbar’s number of 150 meaningful relationships per ego, thus fostering a significantly enriched social life within the Metaverse – an advancement not observed in online relationships within the social web (Dunbar et al., 2015).
Figure 2. Spare time for an ego network with $|V|=117,\tilde{X}=1288h,Z=300h,\beta=1.29$.
6. Conclusion and discussion
@@
This work’s primary contribution lies in addressing social life within the Metaverse through an approach rooted in social sciences findings. It introduces an innovative model that delineates user-mediated and avatar-mediated interactions, synthesizing them into a coherent social cognitive involvement framework. Our focus has been on enhancing user socialization by streamlining interactions through avatar delegation, serving as a foundational step for a more comprehensive model of social interactions in the Metaverse. Immediate advancements could involve exploring various dependencies among problem variables, surpassing mere linearity. Expanding the scope, one intriguing direction involves integrating the concept of social circles present in Dunbar’s ego network models, an aspect overlooked in our current paper. Additionally, a promising avenue for research involves determining the minimum level of user-mediated interactions necessary to maintain relationships within the ego network. Building upon this, we aim to explore how users might reinvest the time saved by avatars into forging new social connections and whether the assistance of avatars leads to a restructuring of user ego networks.
Acknowledgements.
@@
This work was partially supported by SoBigData.it which receives funding from European Union – NextGenerationEU – National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) – Project: “SoBigData.it – Strengthening the Italian RI for Social Mining and Big Data Analytics” – Prot. IR0000013 – Avviso n. 3264 del 28/12/2021. Chiara Boldrini was also supported by PNRR - M4C2 - Investimento 1.4, Centro Nazionale CN00000013 - “ICSC – National Centre for HPC, Big Data and Quantum Computing” – Spoke 6, funded by the European Commission under the NextGeneration EU programme. Andrea Passarella was partly supported by PNRR – M4C2 – Investimento 1.3, Partenariato Esteso PE00000013 – “FAIR - Future Artificial Intelligence Research” – Spoke 1 “Human-centered AI”, funded by the European Commission under the NextGeneration EU programme.

### v2

@@ III-C Social presence
In the social psychology literature, our notion of “independent avatar” would naturally fall in the category of “agent”. Research works which have investigated the impact of the (perceived) agency on social presence typically assume the user as an actual person or a computerized character prior to the interaction. As the survey [17] nicely points out, the majority of the related literature converges to the conclusion that avatars generally elicit greater social presence value than agents, in the sense that users typically feel higher levels of social presence when a virtual counterpart was thought to be controlled by an actual person rather than by a computer program.
@@
$y_{v}=f_{s}(\tilde{x}_{v})=\beta_{v}\cdot\tilde{x}_{v}$ (8)
time, with, according to [33],
$\mathbf{E}[\beta_{v}]=\frac{4.182}{3.236}=1.29,$ (9)
where $\tilde{x}_{v}$ would be the time that $u$ would need to spend socializing with $v$ instead. $4.182$ is the average social presence value measured from human-human interactions in [33], while $3.236$ is the average social presence value measured in the same experiment for human-computer interactions.
III-D Social cues and compression
@@
$z_{v}=c\cdot\delta\cdot y_{v}=\gamma\cdot y_{v},$ (11)
@@
According to known works on compression ratios, such as [34] and[35], we can set
$\mathbb{E}[c]=0.54.$ (12)
Also, social presence can be measured in several directions (verbal, visual, auditory, etc.). [18] provides a quantitative survey-based assessment that captures both the visual and verbal efficiency of chatbots, using a 7-point Likert scale. As an example, in the best case, highly anthropomorphic chatbots with a human-like conversation style reach a score of 6 out of 7. We use these results to assign a value to $\delta$. Specifically, in the above example, we say that
$\mathbb{E}[\delta]=7/6.$ (13)
@@
IV Maximizing the user spare time
@@ IV-A Problem formulation
For a user $u$ and independent avatar $a$, given a set of contacts $V$ and their respective initial non-avatar-mediated time partitioning $\tilde{x_{1}}+...+\tilde{x_{n}}=\tilde{X}$, maximize the spare time of $u$ (i.e., $\tilde{X}-X-Z$, where $X$ denotes the non-avatar-mediated time for user $u$ when using an avatar to partially offload social interactions) by also using $a$’s avatar-mediated time for socialization in a resulting total combined time partitioning $(x_{1}+y_{1})+...+(x_{n}+y_{n})=X+Y$.
@@
$\min\quad$ $\sum_{v}x_{v}+\gamma\cdot y_{v}$ (14)
@@
$x_{v}+\frac{1}{\beta_{v}}\cdot y_{v}\geq\tilde{x}_{v}$ $\forall v$ (15)
$\sum_{v}x_{v}+\gamma\cdot y_{v}\leq\tilde{X}$ (16)
$\sum_{v}y_{v}\leq Y$ (17)
$\sum_{v}\gamma\cdot y_{v}\leq Z$ (18)
$x_{v},y_{v}\geq 0$ $\forall v$ (19)
The objective function (Eq. 14) minimizes the sum of the non-avatar-mediated time and the debriefing time, thus maximizing the available non-avatar-mediated spare time on the user side. Constraints in Eq. 15 guarantee that the resulting effective non-avatar-mediated and avatar-mediated socializing time for each $v$ is not less than the initial non-avatar-mediated time, i.e., that the resulting relationships are not modified in terms of cognitive involvement with respect to the case when the avatar is not present. Constraint in Eq. 16 guarantees that the non-avatar-mediated time and the debriefing time do not surpass the total available time of the user. Constraint in Eq. 17 guarantees that the avatar-mediated time does not surpass the total available time of the avatar. Constraint 18 guarantees that the debriefing process abides by the non-distractibility rule introduced with Eq. 6. Finally, constraints in Eq. 18 guarantee that time values are zero or positive.
IV-B Problem feasibility
@@
$x_{v}=\tilde{x}_{v}-\frac{1}{\beta}y_{v}.$ (20)
Effectively, we just need to find a solution for $y_{v}$ and the values for $x_{v}$ will follow. This allows us to rewrite the optimization problem as shown below.
$\min\quad$ $\tilde{X}+\left(\gamma-\frac{1}{\beta}\right)\sum_{v}y_{v}$ (21)
@@
$\left(\gamma-\frac{1}{\beta}\right)\sum_{v}y_{v}\leq 0$ (22)
$\sum_{v}y_{v}\leq\min\left\{Y,\frac{Z}{\gamma}\right\}$ (23)
$0\leq y_{v}\leq\beta\tilde{x}_{v}$ $\forall v$ (24)
[…]
@@
V Insights from the model
[…]
@@ V Insights from the model
To quantify the potential impact of the spare time, we compare the saved time against the time an average user spends on individual alters of its ego network layers. We extract these numbers from [13], where statistics for three concentric layers are reported. The layers are referred to as the “support clique”, “sympathy group”, and “active network”, with average cumulative sizes of 4.6, 14.3, and 132.5 alters, respectively. This structure deviates from the standard ego network structure due to the exclusion of the “affinity group” layer, whose properties are not well-defined in the related literature. From [13], we obtain $t_{\textrm{active}}=8.81h$, $t_{\textrm{sympathy}}=38.72h$, $t_{\textrm{support}}=74h$. Considering together the time spent on all the alters in each layer, we obtain $T_{\textrm{active}}=1041.51h$, $T_{\textrm{sympathy}}=375.83h$, $T_{\textrm{support}}=340.65h$. It is evident that when the amount of time saved is high enough, users might have the opportunity to reinvest it in expanding their ego network circles and enhancing their social connections. For instance, with approximately 1000 spare hours, one could potentially double the active layer of their ego network, or with around 350 spare hours, strengthen their support clique. While fully characterizing the dynamics of this augmentation is beyond the scope of this work, and the relationship between time saved and new connections gained might not be as direct and straightforward, these initial insights suggest that avatars in the Metaverse have the potential to increase a user’s available time – a resource that could be redirected toward growing their personal social network. In theory, this could result in surpassing Dunbar’s number of 150 meaningful relationships per ego, thus fostering a significantly enriched social life within the Metaverse – an advancement not observed in online relationships within the social web [16].
Figure 4: Spare time for an ego network with $|V|=117,\tilde{X}=1288h,Z=300h,\beta=1.29$.
VI Conclusion and discussion
This work’s primary contribution lies in addressing structural properties of human social networks in the Metaverse, through an approach rooted in social sciences findings. We posit that this is a fundamental building block for the design of Metaverse social networks, as similar models have been for the design of online social networks. Leveraging those concepts, our approach is as follows:
  1. 1.
    Navigate the network of personal connections within the Metaverse through the anthropological lens of ego networks, focusing on the user related parameter configuration and the relation between physical and digital interactions.
  2. 2.
    Incorporate social presence concepts to allows us to delve into the perceptual aspects of these digital relationships, emphasizing the sense of being together in the virtual space and the impact of this presence on users’ social experiences.
  3. 3.
    Integrate insights from social cueing observations, so as to quantify cues embedded in digital interactions to enhance the authenticity and richness of social experience modelling.
The paper introduces an innovative model that delineates user-mediated and avatar-mediated interactions, synthesizing them into a coherent social cognitive involvement framework. Our focus has been on enhancing user socialization by streamlining interactions through avatar delegation, serving as a foundational step for a more comprehensive model of the characteristics of users’ social networks in the Metaverse.
@@ VI Conclusion and discussion
Immediate advancements could involve exploring various dependencies among problem variables, surpassing mere linearity. Expanding the scope, one intriguing direction involves integrating the concept of social circles present in Dunbar’s ego network models, an aspect overlooked in our current paper. Additionally, a promising research direction involves determining the minimum level of user-mediated interactions necessary to maintain relationships within the ego network. Building upon this, we aim to explore how users might reinvest the time saved by avatars into forging new social connections and whether the assistance of avatars leads to a restructuring of user ego networks.
Acknowledgments
@@
This work was partially supported by SoBigData.it, which receives funding from the European Union – NextGenerationEU – National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) – Project: “SoBigData.it – Strengthening the Italian RI for Social Mining and Big Data Analytics” – Prot. IR0000013 – Avviso n. 3264 del 28/12/2021. Chiara Boldrini was supported by PNRR - M4C2 - Investimento 1.4, Centro Nazionale CN00000013 - “ICSC – National Centre for HPC, Big Data and Quantum Computing” – Spoke 6, funded by the European Commission under the NextGeneration EU programme. Marco Conti and Andrea Passarella were partly supported by PNRR – M4C2 – Investimento 1.3, Partenariato Esteso PE00000013 – “FAIR - Future Artificial Intelligence Research” – Spoke 1 “Human-centered AI”, funded by the European Commission under the NextGeneration EU programme.